OpenStudy (anonymous):

Word Problem: If one side of a right triangle is 3 inches and the hypotenuse is 4 inches, how long is the other side of the triangle?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Egyptian triangle

OpenStudy (anonymous):

Hypotenuse is 4 or 5?

OpenStudy (anonymous):

sqart7 i think

OpenStudy (anonymous):

Hypotenus is 4 inches

OpenStudy (anonymous):

\[\sqrt{3^2+4^2}\]=5

OpenStudy (anonymous):

No Victor.

OpenStudy (anonymous):

Isn't it \[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

\[\sqrt{4^{2}-3^{2}}\]

OpenStudy (anonymous):

in other words: \[3^{2}+b ^{2}=4^{2}\]

OpenStudy (anonymous):

sorry, yes it is square root from (4^2-3^2)

OpenStudy (anonymous):

so then it would be: 16-9=7 or \[4^{2}-3^{2}=2.645751311^{2}\]

OpenStudy (anonymous):

Making the other side of the triangle 2.645751311 inches.